OpenStudy (anash):

Ellen wishes to mix candy worth $2.02 per pound with candy $3.31 per pound to form 23 pounds of a mixture worth $2.30 per pound. How many pounds of the more expensive candy should she use?

OpenStudy (anonymous):

3.31 * 7 = 23.17 3.31 * 6 = 19.86 Is it a multiple choice question? @anash

hero (hero):

x = pounds of 1st type of candy y = pounds of 2nd type of candy x + y = 23 2.02x + 3.31y = 23(2.30)